OpenStudy (anonymous):

Complete these ordered pairs for this equation. (-1, ), (0, ), (1, )

OpenStudy (kewlgeek555):

Hm. Which equation? Do you have the slope-intercept form equation? What's the slope? What's the y-intercept?

OpenStudy (anonymous):

2x + y = 3

OpenStudy (kewlgeek555):

Okay. Hm. Let me put this into the slope-intercept form:\[y=-2x+3\]So we knoe that the slope is -2, meaning a negative slope, and we know the y-intercept is 3. So...we have already solved one of these ordered pairs! (0, 3) Now we have to solve for (-1, ___) and (1, ___) To do this we plug in the x-value into the x-variable in the slope-intercept. So, let's solve for the -1 first. \[y=-2(-1)+3\]Can you solve this?

OpenStudy (anonymous):

i dont know how to do any of it.. :(

undeadknight26 (undeadknight26):

cool beans! (thats my term :D) y=−2(−1)+3 y = 2 + 3

undeadknight26 (undeadknight26):

Thats what it would look like right kewlgeek?

OpenStudy (kewlgeek555):

Okay. Look. The (-1) is just like saying "times -1" or "* -1". So let me re-write it.\[y=-2\times-1+3\]Remember PEMDAS. So based on PEMDAS, e should do the multiplication first. I am going to show you a trick with multiplying and dividing integers with the same sign: "SAME SIGNS, POSITIVE ANSWER" So, based on this we know that -2 times -1 will be a positive answer, so let's just do 1 times 2. So you know that 1 times 2 is...

OpenStudy (kewlgeek555):

After you find the answer, then add it by three, then you have the y-value for -1.

OpenStudy (kewlgeek555):

@undeadknight26 was kind enough to give you the answer before I did. ;]

undeadknight26 (undeadknight26):

What can u say right the undead does have soft spots too :D

OpenStudy (kewlgeek555):

So, if you understood what @undeadknight26 and I did, then you should understand what to do to find (1, ___). Do you think you could do it by yourself, do you understand @coolbeans95